One number is 4 less than another. The sum of the numbers is 20. Find the numbers.
Answer by rm29924(97)

let x= a number
one number is 4 less which means you subtract 4
x-4 = the other number
their sum is 20
so add them together
x+x-4=20
add 4 to both sides
x+x=24
2x=24
divide by 2 to get the variable x by itself
x=12
the other number is x-4
12-4
8
answer: the two numbers are 12 and 8
check your answer by plugging it into the original equation